40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Sep 19 02:33:34 2011, kiwamu, HowTo, General, Plan for this week 
    Reply  Mon Sep 19 11:49:32 2011, Jenne, HowTo, General, Plan for this week: SUS inversion 
    Reply  Mon Sep 19 14:57:36 2011, kiwamu, Summary, IOO, IP POS is back wiring.png
Message ID: 5455     Entry time: Mon Sep 19 02:33:34 2011     Reply to this: 5456   5459
Author: kiwamu 
Type: HowTo 
Category: General 
Subject: Plan for this week 

GOAL1:  Stable lock of DRMI

GOAL2:  Measurement of the LSC input matrix in the DRMI configuration

 

 /- - Daytime works - - /

  + Measurement of the arm lengths (Jenne / Kiwamu / volunteers)

  + Optimization of the oplev control loops (Paul)

  + Inversion and installation of the SUS input matrices (Jenne)

  + Tuning of the SUS damping gains (Steve)

  + Measurement of the modulation depths (Mirko)

  + Preparation of the green broadband PD (Katrin)

  + Fixing the Y arm green lock servo (Katrin / Kiwamu)

  + Installation of RFPDs (Anamaria)

  + Minimization of the AM sidebands (Anamaria / Keiko)

  + Preparation of a script for measuring the LSC input matrix (Keiko)

  + MC WFS (Suresh)

  + Online adaptive filtering (Mirko / Jenne)

  + Modification of C1ASS (Kiwamu)

  + Fixing IPPOS (volunteers) 

  + Auto alignment of PRCL and SRCL (volunteers)

  + Loss measurement of the arm cavities (volunteers)

  + Fixing the ETMX SIDE slow monitor (volunteers)

 

 /- - Nighttime works - - /

 + Locking of DRMI

 + Characterization of DRMI and complete the wiki page

ELOG V3.1.3-